Grade 11Magical Mathematics[Interesting Approach]

101^100 -1 is divisible by…

doppalapudi sairam
15 Years agoGrade 11

1 Answer

Chetan Mandayam Nayakar
15 Years ago

101^100 -1=(100+1)^100 -1, is divisible by at least 100C1*100=10000